OpenStudy (anonymous):

Solve the equation: 24x = 8x + 1

OpenStudy (anonymous):

you would minus the 8x n get 16x=1 divide the 16 and get 1/16

OpenStudy (imstuck):

you need to get the x terms on one side of the equals sign together. When you move the 8x over it will be subtracted. 24x - 8x = 1. Can you solve that?

OpenStudy (bradely):

24x=8x+1 24x-8x=1 16x=1 x=1/16 Source:www.mathskey.com

OpenStudy (anonymous):

my final answer i got 1/4 cause u ave to simply

OpenStudy (bradely):

it is wrong

OpenStudy (bradely):

Please check ur question

OpenStudy (anonymous):

Solve the equation: 24x = 8x + 1 Choose one answer. a. 1/4 b. 1/3 c. 3 d. 5

OpenStudy (bradely):

Again please check ur question . options are wrong

OpenStudy (anonymous):

those are the only options on did u simply